OpenStudy (anonymous):

any one knows how to integrate e^-2t cost dt ?

OpenStudy (amistre64):

by parts i believe

OpenStudy (amistre64):

you end up getting the same integral after a few rounds that you can use to your advantage

OpenStudy (amistre64):

\[\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-\int 2e^{-2t}sin(t) dt\] \[\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-2\int e^{-2t}sin(t) dt\] \[\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-2(e^{-2t}cos(t)-\int e^{-2t}cos(t) dt)\]

OpenStudy (amistre64):

hopefully i kept that straight

OpenStudy (amistre64):

nope, i see an error on the last one

OpenStudy (anonymous):

hmm sorry to interrupt but which 1 u use as u and which 1 u use as v for intrating by parts? u = exponential and v= trigo?

OpenStudy (amistre64):

\[\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-2\left( e^{-2t}cos(t)-\int -2e^{-2t}cos(t) dt \right)\] \[\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-2\left( e^{-2t}cos(t)+2\int e^{-2t}cos(t) dt \right)\] \[\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-2e^{-2t}cos(t)-4\int e^{-2t}cos(t) dt\] \[4\int e^{-2t}cos(t) dt+\int e^{-2t}cos(t)dt= -e^{-2t}sin(t)-2e^{-2t}cos(t)\] \[5\int e^{-2t}cos(t) dt= -e^{-2t}sin(t)-2e^{-2t}cos(t)\] \[\int e^{-2t}cos(t) dt= \frac{-e^{-2t}sin(t)-2e^{-2t}cos(t)}{5}\] maybe??

OpenStudy (amistre64):

it shouldnt matter which one you use since they both cycle

OpenStudy (amistre64):

I used the e^ to go down, and the cos to go up

OpenStudy (amistre64):

......at least thats what I had in mind .... might of got things crossed in my head tho ;)

OpenStudy (anonymous):

integrate by parts

OpenStudy (anonymous):

thx dude i will read that out million thx man

OpenStudy (amistre64):

youre welcome; if anything, redo it and go thru 2 integration by parts; then you can use the process of swapping sides and such

OpenStudy (anonymous):

ok thx :)

OpenStudy (amistre64):

yeah, I see what I did. I inadvertently differentiated both of them instead so my answer is off on that regards; but the concept is valid lol

OpenStudy (anonymous):

u mean u differentiated both instead of differentiating 1 and integrating 1?

OpenStudy (amistre64):

yeah, trying to work in in my head i got a few things crossed; let me try it again correctly tho

OpenStudy (anonymous):

okok i will do it too hope i can get it right..

OpenStudy (amistre64):

\begin{array}l &&cos(t)\\ +&e^{-2t}&sin(t)\\ -&-2e^{-2t}&-cos(t)\\ +&4e^{-2t} \end{array} \[\int e^{-2t}cos(t)dt= e^{-2t}sin(t)-\int-2e^{-2t}sin(t)\ dt\] \[\int e^{-2t}cos(t)dt= e^{-2t}sin(t)+2\int e^{-2t}sin(t)\ dt\] \[\int e^{-2t}cos(t)dt= e^{-2t}sin(t)+2\left(-e^{-2t}cos(t)-\int 2e^{-2t}cos(t)\ dt\right)\] \[\int e^{-2t}cos(t)dt= e^{-2t}sin(t)-2e^{-2t}cos(t)-4\int e^{-2t}cos(t)\ dt\] now we add 4(int: e^(-2t) cos(t)) to both sides. \[4\int e^{-2t}cos(t)\ dt+\int e^{-2t}cos(t)dt= e^{-2t}sin(t)-2e^{-2t}cos(t)\]combine like terms \[5\int e^{-2t}cos(t)\ dt= e^{-2t}sin(t)-2e^{-2t}cos(t)\]and divide of the 5 \[\int e^{-2t}cos(t)\ dt=\frac{ e^{-2t}sin(t)-2e^{-2t}cos(t)}{5}\] and then we dbl chk with the wolf ;)

OpenStudy (anonymous):

hey we got the same answer yess but it has a constant rite? do i need to add a + K?

OpenStudy (amistre64):

yes, the +C at the end is good to put there since this has no definite result

OpenStudy (anonymous):

yes okk man thxx !